OpenStudy (anonymous):

Geraldo’s current salary is $40,000 per year. His annual pay raise is always a percent of his salary at the time. What would his salary be if he got four consecutive 4% increases? a. $46,794.34 c. $43,328.04 b. $52,697.23 d. $62,156.12

OpenStudy (cruffo):

You can solve this problem using a series of calculations New Salary = Old Salary + %increase * Old Salary Recall that 4% = 0.04 as a decimal number Salary0 = 40,000 Salary1 = 40,000 + 0.04(40,000) <- First increase Salary2 = Salary1 + 0.04(Salary1) <- second increase Salary3 = Salary2 + 0.04(Salary2) <- third increase Salary4 = Salary3 + 0.04(Salary3) <- forth increse

OpenStudy (cruffo):

This is basically the same as annual compound interest A = P(1+r)^t $40,000 at 4% compounded annually for 4 years. P = 40000 r = 0.04 t = 4 A = 40000(1 + 0.04)^4